# Can language models learn analogical reasoning? Investigating training objectives and comparisons to human performance

### v1

@@ 3.3.5 External Tasks: Semantic Similarity
In order to see if our training scheme affects performance on external tasks, as can be the case with catastrophic forgetting, we test our models on three word-level, non-contextual semantic similarity datasets: SimLex-999, MEN, and WordSim353 (WS353) Hill et al. (2015); Bruni et al. (2014); Finkelstein et al. (2001); Kemker et al. (2018)777 We used some code from https://github.com/kudkudak/word-embeddings-benchmarks. All these datasets contain words pairs with a similarity measure, however Hill et al. (2015) details some key features of how these dataset differ when they introduced SimLex-999; namely that both the MEN and WS353 tended to measure word relatedness/association as opposed to word similarity (not that these are mutually exclusive), and MEN’s tendency to focus on less abstract concepts, such as nouns. The similarity measure within these datasets range from 0 to 10, while we use cosine similarity (details described in the next section), which gives a similarity measure in the range -1 to 1. We chose these tasks because it is a word level similarity task, which is related to our analogy task. Ideally the performance on these tasks would improve with our training, or minimally not decrease.
Table 3: Average # of times entities seen in pre-training data by true label and model guess before/after fine tuning (BERT a-b)
@@
True Positive 205376 108667 187234 181359 185524
True Negative 207319 141762 170619 199694 191815
206328 126385 182640 194465
@@
Table 5: Average Accuracy on Distractor Dataset. Arrows next to BERT a-b and Single Layer a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
@@
Simple BERT a-b Single Layer a-b BERT a-c
Distractor Salience Distractor Salience Distractor Salience Distractor Salience
Overall High Low Overall High Low Overall High Low Overall High Low
OVERALL 0.52 0.51 0.54 0.70 0.69↑ 0.71↑ 0.72 0.72 0.73 0.43↑ 0.33↑ 0.53↑
Near Overall 0.63 0.56 0.70 0.76 0.72↑ 0.80↑ 0.80 0.77 0.83 0.39↑ 0.31↑ 0.47↑
Categorical 0.64 0.57 0.70 0.76↑ 0.74↑ 0.77↑ 0.81↑ 0.74↑ 0.87↑ 0.44 0.34↑ 0.53
Causal 0.63 0.58 0.67 0.72↑ 0.66↑ 0.77↑ 0.78↑ 0.73↑ 0.83↑ 0.38↑ 0.27↑ 0.48↑
Compositional 0.63 0.53 0.73 0.81 0.77↑ 0.85↑ 0.83 0.85↑ 0.80↑ 0.36↓ 0.32↓ 0.39↓
Far Overall 0.42 0.46 0.37 0.63↑ 0.65↑ 0.61↑ 0.64↑ 0.66↑ 0.63↑ 0.47↑ 0.34↑ 0.60↑
Categorical 0.46 0.55 0.37 0.68↑ 0.74↑ 0.61↓ 0.78↑ 0.83↑ 0.72↑ 0.49↑ 0.32↑ 0.66↑
Causal 0.36 0.38 0.34 0.68↑ 0.69↑ 0.66↑ 0.53↑ 0.56↑ 0.50↑ 0.45↑ 0.40↑ 0.49↑
Compositional 0.44 0.46 0.41 0.55↑ 0.52↑ 0.57↑ 0.63↑ 0.59↑ 0.66↑ 0.48↓ 0.31↑ 0.65↓
@@ 4.1 Proportional Analogies as a Learnable Objective for Neural Networks
Table 1 shows accuracy on classifying the testset with both the baselines and trained models. FastText has a tendency to label all analogies as negative given the cosine similarity measure, while BERT models have a tendency to classify all analogies as positive. This is perhaps unsurprising, as it has been demonstrated that word embeddings exhibit anisotrophy, and that anisotropy is higher with contextual word embeddings, resulting in any two random words having high cosine similarity to each other, perhaps translating into the distances between words being similar to each other Ethayarajh (2019). BERT a-b seems to be less likely to be biased towards one label as compared to the other baselines, with the relatively large SCAN dataset having the greatest tendency to be classified as positive.
Both a-b training schemes improved overall accuracy on analogy classification over the previously discussed baselines, with most positive changes in accuracy being statistically significant. The largest gains were with the SCAN dataset, mostly due to an increased ability to correctly classify negative analogies. Performance was generally better with the science analogies than metaphors, with both a-b models reaching over 0.85 accuracy overall. Czinczoll et al. (2022) found that models performed better on science analogies than on metaphor analogies, which they attributed to metaphors being more abstract. As mentioned before, while the model would have never seen the examples from the evaluation set, it would have seen the entities in the pairs that make up the samples in the evaluation set as parts of other analogies. There was improvement on the other datasets with both a-b models, however the overall improvements were less dramatic, ranging from +0.04 on the U2 data set when comparing Single Layer a-b to untrained BERT a-b to +0.11 on the U4 dataset Single Layer a-b vs. BERT a-b non-tuned. We cannot directly compare our results to Czinczoll et al. (2022) and Ushio et al. (2021b), as Czinczoll et al. (2022) does not use a classification or ranking multiple choice task while Ushio et al. (2021b) used the entire list of negative analogies for the ranking task, and both the main goals of these papers were not to test training schemes.
The fine-tuned and non-fine-tuned models were generally able to perform an analogy ranking task better than the classification task, as shown in Table  2. The performance of each model between SCAN and the other datasets was less variable with the ranking task as opposed to the classification task. Again, the fine-tuned models outperformed the baselines with statistically significant improvements.
@@ 4.2 Exploring Accuracy in Relation to Word Frequency and Subwords
Table 3 shows average word frequency by true and predicted label, while Table 4 shows classification accuracy by whether an analogy had at least one OOV entity. The entities contained in false analogies tended to be observed in the pre-training data more frequently than those in true analogies. However, analogies predicted as being true analogies contained entities that were on average seen a little over 60% more on average than those contained in analogies predicted as false before training. Additionally, analogies that contained no OOV entities were almost always predicted as true before training (Table 4). After training, the average frequency among predicted labels closely matches that among the true labels, and accuracy improved greatly among negative analogies with no OOV entities, as well as among analogies with OOV entities. It appears that before fine-tuning, the model overestimated the similarity in relations between analogy pairs with in-vocabulary words, and a bulk of the learning effected the ability to correctly identify lack of analogy. Similar trends can be seen when looking within each dataset (Tables 9-16 in appendix A).
@@ 4.3 Neural Networks as Compared to Humans
Table 6: Average Performance on Several Semantic Similarity Tasks. Arrows next to BERT a-b and Single Layer a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT non-tuned. BERT a-c is also compared to BERT non-tuned, since each involves just one word so the objective function without training is irrelevant. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
@@
FastText BERT Simple BERT Single Layer BERT
Tasks non-tuned Classification a-b a-b a-c
SimLex-999 0.44 0.17 0.21 0.33 0.34 0.19↑
MEN 0.81 0.27 0.27 0.31 0.28↑ 0.33
WS535 0.69 0.27 0.27 0.27→ 0.26↓ 0.27→
In our experiments, the best performing model was Single Layer a-b, with a 0.72 overall accuracy, up from 0.53 with BERT a-b non-tuned, and $\approx$0.12 worse than human performance. Accuracy for both BERT a-b and Single Layer a-b mostly increased with training over the baseline, however most increases among subgroups were not statistically significant, notably the sample size was small. When looking at subgroups, the same trends observed in humans were not present, nor were there any obvious trends among subgroups between the models, with the exception that near analogies were generally easier to solve than far analogies.
@@ 4.4 Performance on External Tasks
Finally, we tested on an external task to find out whether fine-tuning on the task of analogical reasoning might have a (negative) effect on semantic similarity tasks. Table  6 shows the Spearman’s rank-order correlation coefficient for the three Semantic Similarity tasks. BERT a-b and Single Layer a-b improved over BERT a-b non-tuned, showing that training actually improved performance, even if performance is still overall low compared to FastText. FastText outperformed all transformer models on the external tasks, which is unsurprising. Ethayarajh (2019) found that FastText and embeddings from lower layers of BERT outperformed final layer hidden representations from BERT, although they used the first principal component of the embeddings so the results are not directly comparable. Given the tasks are non-contextual, perhaps the contextual nature of BERT that allows it to perform well on certain tasks hinders it in others. Interestingly BERT a-b and Single Layer a-b performed better on the SimLex-999 task than the other two tasks, unlike the baseline models presented. Hill et al. (2015) had found the SimLex-999 task was harder for neural embeddings to solve than MEN or WS353, which they attributed to these models being better at identifying word association than similarity. However they did not test BERT-like models.
@@ 5 Conclusion, Limitations and Future Work
In this paper, we aimed to move from testing relatively simple analogical relations in pre-trained language models to testing the ability to learn more complex relations used for testing human analogical reasoning, with a tailored training objective. We found that overall, analogies are something that can be learned. We reach an accuracy of 0.72 up from 0.53 while being 0.12 below the human upper bound, on an unseen test set constructed to test human analogical reasoning. Lastly, we find that fine-tuning models with certain training objectives generally does not deteriorate their performance on external, but related tasks. In fact, on some tasks we observed improved accuracy.
@@
One direction for future work is to address limitations of word embeddings or their representation power, such as the anisotrophy that exists among contextual word embeddings Ethayarajh (2019). Perhaps a distance measure between two entities that is not subtraction would be a better way to represent their relation. Additionally, since prior work has suggested that analogies where the entities are close to each other in space are generally easier to solve with the vector offset method, perhaps focusing on incorporating knowledge regarding semantic distance between entities during training would be helpful Rogers et al. (2017). We would also like to explore whether augmenting LM training with analogy learning for other common NLP benchmarks affects performance on these benchmarks.
@@ Acknowledgements
We are grateful to the Swiss National Science Foundation $SNSF,grant205121\_207437:C-LING$ for funding this work. We also thank members of the Idiap NLU-CCL group for helpful discussions and suggestions, the NLP lab at EPFL, and the anonymous reviewers for their fruitful comments and suggestions.
@@ Appendix A Characteristics of our Dataset
Table 7: Example of Analogies in the Datasets used for Training
@@
Table 8: Characteristics of Entities in Datasets
@@
Table 9: SCAN dataset: Average # of times entities seen in pre-training data by true label and model guess before/after fine tuning (BERT a-b)
@@
Table 10: SCAN dataset: Accuracy among analogies with no OOV entities and those with at least one before/after fine tuning (BERT a-b)
@@
Table 11: SAT dataset: Average # of times entities seen in pre-training data by true label and model guess before/after fine tuning (BERT a-b)
@@
Table 12: SAT dataset: Accuracy among analogies with no OOV entities and those with at least one before/after fine tuning (BERT a-b)
@@
Table 13: U2 dataset: Average # of times entities seen in pre-training data by true label and model guess before/after fine tuning (BERT a-b)
@@
Table 14: U2 dataset: Accuracy among analogies with no OOV entities and those with at least one before/after fine tuning (BERT a-b)
@@
Table 15: U4 dataset: Average # of times entities seen in pre-training data by true label and model guess before/after fine tuning (BERT a-b)
@@
Table 16: U4 dataset: Accuracy among analogies with no OOV entities and those with at least one before/after fine tuning (BERT a-b)
@@
Appendix B Results from other BERT architectures
We ran a subset of our experiments with different BERT models: BERT-cased (110M parameters), BERT-large-uncased (340M parameters), and RoBERTa-base (125M parameters). Specifically the models we reproduced were BERT a-b non-tuned, BERT a-b, and Simple Classifier. These were all done with a single split of the data.

### v4

@@ 3.3.5 External Tasks: Semantic Similarity
In order to see if our training scheme affects performance on external tasks, as can be the case with catastrophic forgetting, we test our models on three word-level, non-contextual semantic similarity datasets: SimLex-999, MEN, and WordSim353 (WS353) Goodfellow et al. (2014); Kirkpatrick et al. (2017); Hill et al. (2015); Bruni et al. (2014); Finkelstein et al. (2001); Kemker et al. (2018)777 We used some code from https://github.com/kudkudak/word-embeddings-benchmarks. All these datasets contain words pairs with a similarity measure, however Hill et al. (2015) details some key features of how these dataset differ when they introduced SimLex-999; namely that both the MEN and WS353 tended to measure word relatedness/association as opposed to word similarity (not that these are mutually exclusive), and MEN’s tendency to focus on less abstract concepts, such as nouns. The similarity measure within these datasets range from 0 to 10, while we use cosine similarity (details described in the next section), which gives a similarity measure in the range -1 to 1. We chose these tasks because it is a word level similarity task, which is related to our analogy task. Ideally the performance on these tasks would improve with our training, or minimally not decrease.
@@
True Positive 205376 108669 185999 184357 185524
True Negative 207319 141762 168606 200765 191815
206329 126385 181109 196079
Table 3: Average # of times entities seen in pre-training data by true label and model guess before(B)/after(A) fine tuning (BERT a-b)
@@
Simple BERT a-b BERT a-c
Distractor Salience Distractor Salience Distractor Salience
Overall High Low Overall High Low Overall High Low
OVERALL 0.52 0.51 0.54 0.69 0.68↑ 0.70↑ 0.45↑ 0.35↑ 0.54↑
Near Overall 0.63 0.56 0.70 0.75 0.71↑ 0.79↑ 0.45↑ 0.37↑ 0.52↑
Categorical 0.64 0.57 0.70 0.73↑ 0.71↑ 0.75↑ 0.50↑ 0.44↑ 0.55
Causal 0.63 0.58 0.67 0.72↑ 0.65↑ 0.78↑ 0.48↑ 0.41↑ 0.54↑
Compositional 0.63 0.53 0.73 0.82 0.78↑ 0.85↑ 0.37↑ 0.26↓ 0.47↑
Far Overall 0.42 0.46 0.37 0.62↑ 0.64↑ 0.61↑ 0.45↑ 0.33↑ 0.56↑
Categorical 0.46 0.55 0.37 0.67↑ 0.73↑ 0.60↓ 0.46↑ 0.30↑ 0.62↑
Causal 0.36 0.38 0.34 0.65↑ 0.66↑ 0.63↑ 0.45↑ 0.38↑ 0.51↑
Compositional 0.44 0.46 0.41 0.56↑ 0.53↑ 0.59↑ 0.43↓ 0.32↑ 0.54↓
Table 5: Average Accuracy on Distractor Dataset. Arrows next to BERT a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT a-b non-tuned. BERT a-c is compared to BERT a-c non-tuned. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
@@ 4.1 Proportional Analogies as a Learnable Objective for Neural Networks
Table 1 shows accuracy on classifying the testset with both the baselines and trained models. FastText has a tendency to label all analogies as negative given the cosine similarity measure, while BERT models have a tendency to classify all analogies as positive. This is perhaps unsurprising, as it has been demonstrated that word embeddings exhibit anisotropy, and that anisotropy is higher with contextual word embeddings, resulting in any two random words having high cosine similarity to each other, perhaps translating into the distances between words being similar to each other Ethayarajh (2019). BERT a-b seems to be less likely to be biased towards one label as compared to the other baselines, with the relatively large SCAN dataset having the greatest tendency to be classified as positive.
The a-b training scheme improved overall accuracy on analogy classification over the previously discussed baseline, with most positive changes in accuracy being statistically significant. The largest gains were with the SCAN dataset, mostly due to an increased ability to correctly classify negative analogies. Performance was generally better with the metaphor analogies than science, with the a-b model reaching 0.87 accuracy overall. Czinczoll et al. (2022) found that models performed better on science analogies than on metaphor analogies, which they attributed to metaphors being more abstract. As mentioned before, while the model would have never seen the examples from the evaluation set, it would have seen the entities in the pairs that make up the samples in the evaluation set as parts of other analogies. There was improvement on the other datasets with the a-b model, however the overall improvements were less dramatic, with accuracy +0.07 when compared to BERT a-b non-tuned. We cannot directly compare our results to Czinczoll et al. (2022) and Ushio et al. (2021b), as Czinczoll et al. (2022) does not use a classification or ranking multiple choice task while Ushio et al. (2021b) used the entire list of negative analogies for the ranking task. Moreover, the main goals of these papers were not to test training schemes.
The fine-tuned and non-fine-tuned models were generally able to perform an analogy ranking task better than the classification task, as shown in Table  2. The performance of each model between SCAN and the other datasets was less variable with the ranking task as opposed to the classification task. Again, the fine-tuned models outperformed their respective baselines with statistically significant improvements, the improvements being much greater with the a-b scheme.
@@ 4.2 Exploring Accuracy in Relation to Word Frequency and Subwords
Table 3 shows average word frequency by true and predicted label, while Table 4 shows classification accuracy by whether an analogy had at least one OOV entity. The entities contained in false analogies tended to be observed in the pre-training data more frequently than those in true analogies. However, analogies predicted as being true analogies contained entities that were seen a little over 60% more on average than those contained in analogies predicted as false before training. Additionally, analogies that contained no OOV entities were almost always predicted as true before training (Table 4). After training, the average frequency among predicted labels closely matches that among the true labels, and accuracy improved greatly among negative analogies with no OOV entities, as well as among analogies with OOV entities. It appears that before fine-tuning, the model overestimated the similarity in relations between analogy pairs with in-vocabulary words, and a bulk of the learning affected the ability to correctly identify lack of analogy. Similar trends can be seen when looking within each dataset.
@@ 4.3 Neural Networks as Compared to Humans
FastText BERT Simple BERT BERT
Tasks non-tuned Classification a-b a-c
SimLex-999 0.44 0.17 0.21 0.33 0.19↑
MEN 0.81 0.27 0.27 0.31 0.33
WS535 0.69 0.27 0.27 0.27→ 0.30↑
Table 6: Average Performance on Several Semantic Similarity Tasks. Arrows next to BERT a-b are labeling whether accuracy went up↑, down↓, or stayed the same→, as compared to BERT non-tuned. BERT a-c is also compared to BERT non-tuned, since this task compares the similarity between two words, not between the differences between two words. Boldface indicates a statistically significant change (p<0.05) using a z-test for proportions.
In our experiments, the best performing model was BERT a-b, with a 0.69 overall accuracy, up from 0.53 with BERT a-b non-tuned, and $\approx$0.15 worse than human performance. Accuracy for BERT a-b mostly increased with training over the baseline, however most increases among subgroups were not statistically significant, although notably the sample size was small. When looking at subgroups, the same trends observed in humans were not present, nor were there any obvious trends among subgroups between the models, with the exception that near analogies were easier to solve than far analogies for our best model.
@@ 4.4 Performance on External Tasks
Finally, we tested on an external task to find out whether fine-tuning on the task of analogical reasoning might have a (negative) effect on semantic similarity tasks. Table  6 shows the Spearman’s rank-order correlation coefficient for the three Semantic Similarity tasks. BERT a-b improved over BERT non-tuned, showing that training actually improved performance, even if performance is still overall low compared to FastText. FastText outperformed all transformer models on the external tasks, which is unsurprising. Ethayarajh (2019) found that FastText and embeddings from lower layers of BERT outperformed final layer hidden representations from BERT, although they used the first principal component of the embeddings so the results are not directly comparable. Given the tasks are non-contextual, perhaps the contextual nature of BERT that allows it to perform well on certain tasks hinders it in others. Interestingly BERT a-b performed better on the SimLex-999 task than the other two tasks, unlike the baseline models presented. Hill et al. (2015) had found the SimLex-999 task was harder for neural embeddings to solve than MEN or WS353, which they attributed to these models being better at identifying word association than similarity. However they did not test BERT-like models.
@@ 5 Conclusion, Limitations and Future Work
In this paper, we aimed to move from testing relatively simple analogical relations in pretrained language models to testing the ability to learn more complex relations used for testing human analogical reasoning, with a tailored training objective. We found that overall, analogies are something that can be learned. We reach an accuracy of 0.69 up from 0.53 while being 0.15 below the human upper bound, on an unseen test set constructed to test human analogical reasoning. Lastly, we find that fine-tuning models with certain training objectives generally does not deteriorate their performance on external, but related tasks. In fact, on some tasks we observed improved accuracy.
@@
One direction for future work is to address limitations of word embeddings or their representation power, such as the anisotropy that exists among contextual word embeddings Ethayarajh (2019). Perhaps a distance measure between two entities that is not subtraction would be a better way to represent their relation. Additionally, since prior work has suggested that analogies where the entities are close to each other in space are generally easier to solve with the vector offset method, perhaps focusing on incorporating knowledge regarding semantic distance between entities during training would be helpful Rogers et al. (2017). We would also like to explore whether augmenting LM training with analogy learning for other common NLP benchmarks affects performance on these benchmarks.
@@ Acknowledgements
We are grateful to the Swiss National Science Foundation ($grant~{}205121\_207437:C-LING$) for funding this work. We also thank members of the Idiap NLU-CCL group, the NLP lab at EPFL, the Programme group Psychological Methods of the University of Amsterdam for helpful discussions, and the anonymous reviewers for their fruitful comments and suggestions.
@@ Appendix A Characteristics of our Dataset
Tables 7 and 8 as well as Figure 1 characterize the training data that we used to reach a better understanding of what the model actually ’sees’ during training. For explanations on how we created the training data from the original datasets we refer the reader to the section on Datasets. Table 7 gives examples of analogies and their negative counterparts used in the data. Table 8 shows some characteristics of the entities that make up each dataset we used for training our models.
@@
Table 7: Example of Analogies in the Datasets used for Training
@@
Appendix B Accuracy in Relation to Word Frequency and Subwords: Within Datasets
Tables 9 - 16 reproduce tables 3 and 4 in the manuscript but for each dataset individually. Similar trends seen in the overall data are seen within each dataset, specifically the improvements among being able to classify false analogies correctly that contain no OOV entities, as well as the model becoming less biased with regards to the number of times entities were seen and the label it chose.
Table 9: SCAN dataset: Average # of times entities seen in pre-training data by true label and model guess before(B)/after(A) fine tuning (BERT a-b)
@@ Appendix B Accuracy in Relation to Word Frequency and Subwords: Within Datasets
Table 10: SCAN dataset: Accuracy among analogies with no OOV entities and those with at least one before(A)/after(B) fine tuning (BERT a-b)
@@
Table 11: SAT dataset: Average # of times entities seen in pre-training data by true label and model guess before(A)/after(B) fine tuning (BERT a-b)
@@
Table 12: SAT dataset: Accuracy among analogies with no OOV entities and those with at least one before(A)/after(B) fine tuning (BERT a-b)
@@
Table 13: U2 dataset: Average # of times entities seen in pre-training data by true label and model guess before(A)/after(B) fine tuning (BERT a-b)
@@
Table 14: U2 dataset: Accuracy among analogies with no OOV entities and those with at least one before(A)/after(B) fine tuning (BERT a-b)
@@
Table 15: U4 dataset: Average # of times entities seen in pre-training data by true label and model guess before(A)/after(B) fine tuning (BERT a-b)
@@
Table 16: U4 dataset: Accuracy among analogies with no OOV entities and those with at least one before(A)/after(B) fine tuning (BERT a-b)
@@
Appendix C Results from architectures: Single layer
We tested a model where we added a single feed forward layer on top of a pretrained BERT model, before the pooling layer. We kept the pretrained BERT weights frozen and trained only the additional layer using the a-b training scheme. Training with a single layer saw minimal change over the BERT a-b non-tuned baseline presented in the main text.
Table 17: Single Layer BERT a-b: Accuracy on Analogy Classification and Ranking Task
Classification Ranking
Category Overall Positive Negative
OVERALL 0.53 0.72 0.33 0.69
SAT 0.53 0.61 0.45 0.63
U2 0.52 0.63 0.40 0.72
U4 0.48 0.62 0.34 0.70
SCAN 0.54 0.79 0.28 0.70
SCAN -Science 0.54 0.76 0.31 0.72
SCAN -Metaphor 0.54 0.79 0.28 0.70
Table 18: Single Layer BERT a-b: Accuracy on Distractor Dataset
Semantic Relation Distractor Salience
Distance Type Overall High Low
OVERALL 0.54 0.52 0.56
Near Overall 0.55 0.53 0.57
Categorical 0.57 0.51 0.62
Causal 0.55 0.55 0.55
Compositional 0.54 0.53 0.55
Far Overall 0.53 0.51 0.55
Categorical 0.66 0.6 0.72
Causal 0.44 0.41 0.46
Compositional 0.49 0.51 0.46
Appendix D Results from other architectures: Different Models
We ran a subset of our experiments with different BERT models: BERT-cased (110M parameters), BERT-large-uncased (340M parameters), and RoBERTa-base (125M parameters). Specifically the models we reproduced were BERT a-b non-tuned, BERT a-b, and Simple Classifier.
On the classification task, all the models saw some increase in overall accuracy as compared to the baseline, with RoBERTa achieving the highest overall accuracy (Tables 19-21). RoBERTa slightly outperformed BERT-base on the classification and ranking tasks (Table 21). The untrained RoBERTa a-b started out with a more extreme difference in accuracy between positive and negative samples in the classification task as compared to the BERT a-b, and while training closed the gap between the accuracy (while improving overall accuracy), BERT a-b was closer to achieving parity between the two groups. RoBERTa’s performance did not transfer as well to the Distractor Dataset as compared to BERT a-b (Table 27). BERT-large-uncased did not see as much improvement with fine-tuning, likely due to the model being too large to learn from the relatively small dataset (Table 20,23,26).